OpenStudy (anonymous):

Graph the function, then answer the questions about the graph. (please show work and explain!!! thanks!) y=1/2x^2-x-6 Vertex:________ x-intercepts:_________ y-intercept:_________ ***and GRAPH!! :D thanks!! :)

OpenStudy (anonymous):

4 parts: 1, find the vertex 2, find the x-int. 3, find the y-int. 4. GRAPH The equation!! :D

OpenStudy (anonymous):

(1/2x^2)-x-6?

jimthompson5910 (jim_thompson5910):

first identify a, b, and c

OpenStudy (anonymous):

a=1/2 b=-1 c=-6 ???

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

if we have an equation in standard form y = ax^2+bx+c, we can easily spot the y-intercept

jimthompson5910 (jim_thompson5910):

plug in x = 0 to get y = ax^2+bx+c y = a(0)^2+b(0)+c y = 0a + 0b + c y = c So when x = 0, y = c This means that the y-intercept for y = ax^2+bx+c is the point (0, c)

jimthompson5910 (jim_thompson5910):

So that part is relatively easy (compared to the other parts)

OpenStudy (anonymous):

so what?? y-intercept is (0,-6) ??

jimthompson5910 (jim_thompson5910):

yes, you replace c with -6

OpenStudy (anonymous):

how do i find the other stuff tho?

jimthompson5910 (jim_thompson5910):

The steps I showed were in general. So if you see y = ax^2+bx+c, you can jump to the y-intercept of (0, c) This means that if you see y = 1/2x - x - 6, you can jump to the y-intercept of (0, -6) quickly

OpenStudy (anonymous):

oh okay i see :) what about x-int? how do i find that for my problem?

jimthompson5910 (jim_thompson5910):

Do you remember how to find the vertex?

jimthompson5910 (jim_thompson5910):

To find the x-intercepts, plug in y = 0 and solve for x

OpenStudy (anonymous):

(h,k)=vertex right?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

how am i finding the x-intercepts??

OpenStudy (anonymous):

can u pls show me?? I'm a little confused :(

jimthompson5910 (jim_thompson5910):

Finding x-intercepts. Plug in y = 0 and solve for x y=1/2x^2-x-6 0=1/2x^2-x-6 2*0 = 2*1/2x^2 - 2*1x - 2*6 0 = x^2 - 2x - 12 x^2 - 2x - 12 = 0

jimthompson5910 (jim_thompson5910):

Now solve x^2 - 2x - 12 = 0 for x

OpenStudy (anonymous):

so a=1 b=-2 c=-12 ?? what do i do with that?

jimthompson5910 (jim_thompson5910):

plug them into the quadratic formula

jimthompson5910 (jim_thompson5910):

\[\Large x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):